OpenStudy (anonymous):

Find the arc length of a semicircle where the diameter is 19 yards. A. 19π/2 B. 19π/4 C. 19π/6 D. 19π/8

OpenStudy (anonymous):

Arc length of a circle is given as \[\huge l = \frac{ \theta \times 2\pi r }{360^0}\] In case of semicircle \[\huge \theta = 180^0\] hence arc length of semi circle =\[\huge = \frac{1}{2} \times \frac{ 180 ^0 \times 2\pi \times 19 }{360^0} =\frac{19 \pi }{2}\] Thus option A. 19π/2 is the correct answer... @amberk123